# Euclidean Distance Matrix Completion via Asymmetric Projected Gradient Descent

### v1

@@ Abstract
This paper proposes and analyzes a gradient-type algorithm based on Burer-Monteiro factorization, called the Asymmetric Projected Gradient Descent (APGD), for reconstructing the point set configuration from partial Euclidean distance measurements, known as the Euclidean Distance Matrix Completion (EDMC) problem. By paralleling the incoherence matrix completion framework, we show for the first time that global convergence guarantee with exact recovery of this routine can be established given $\mathcal{O}(\mu^{2}r^{3}\kappa^{2}n\log n)$ Bernoulli random observations without any sample splitting. Unlike leveraging the tangent space Restricted Isometry Property (RIP) and local curvature of the low-rank embedding manifold in some very recent works, our proof provides new upper bounds to replace the random graph lemma under EDMC setting. The APGD works surprisingly well and numerical experiments demonstrate exact linear convergence behavior in rich-sample regions yet deteriorates fast when compared with the performance obtained by optimizing the s-stress function, i.e., the standard but unexplained non-convex approach for EDMC, if the sample size is limited. While virtually matching our theoretical prediction, this unusual phenomenon might indicate that: (i) the power of implicit regularization is weakened when specified in the APGD case; (ii) the stabilization of such new gradient direction requires substantially more samples than the information-theoretic limit would suggest.
Index Terms:
Euclidean distance matrix completion, Burer-Monteiro factorization, Dual Basis.
@@ I Introduction
Given $n$ points $\{\mathbf{p}_{i}\}_{i=1}^{n}$ embedded in $\mathbb{R}^{r}$, $r=2,3$, calculating their inter-point distances is a trivial task. Forwarding the perfect, complete $n(n-1)/2$ distances back to the point set configuration is, meanwhile well-known as the classic Multi-dimensional Scaling (cMDS)[2, Sec. 3.1.1][3][4, Thm. 2]. However, significant efforts have been conducted to handle the case when part of the distance measurements are missing, known as the Euclidean Distance Matrix Completion (EDMC) problem[5][4]. This not only dues to its ubiquitous presence in, e.g., computational chemistry[6][7][8][9], operating wireless sensor networks[10][11][12] and articulated robots[5, Sec. 4.3.2][13], channel estimation[14], indoor SLAM[15], ultrasound calibration[16], but also the fact that this inverse problem is NP-hard to solve in general[17, Sec. 3.4]. To escape from the NP-hardness, one may only consider a selected region of the ground turth (generic[18][19], incoherence[20][21]), enforcing certain connectivity of the underlying sample graph (Erdös-Rényi[22][23], random geometric graph[12][24][25][26]), and posing constrains on certain geometry of the point set ($r$-unique localizable[27], trilateration[28], universal rigidity[29]). Therefore, performance guarantee results in this field can be roughly divided into two classes, i.e., combinatorial[27][29][30] and probabilistic. While the latter can be further categorized into the traditional yet effective view from random geometry[12][24][25][26], and the recent revisit of EDMC via incoherent matrix completion lens[20][23][31][22]. Method proposed in this manuscript belongs to the last class. In short, we analyze an algorithm analogous to the Iterative Fast Hard Thresholding for EDMC (IFHT-EDMC) that appeared in a very recent work [23], under standard incoherence assumption[32] and Bernoulli sampling scheme, from the quotient geometry perspective.
I-A Related Work
We are interested in answering the following question
@@ I-A Related Work
Q1: Does the theoretical success of the two-step non-convex factorization type incoherent matrix completion algorithms [33][34][35][36] generalize to the EDMC setting?
@@
While the Bernoulli sample model used in these works does not have strong real-world application scenarios when specified to EDMC, we are not aware of any work that analyzed the non-convex gradient refinement stage under a more practical sample model that depends on the magnitudes of EDM entities, i.e., the so-called unit ball rule, under the random geometry setup [^1] We emphasize that the transition bound obtained from random geometry graph is often sharp up to constant.. Javanmard and Montanari developed the stability notion of a rigid graph[12, Lemma 4.1, 4.2] and use that to show the perturbation on convex solution can be bounded. [26][24] analyzed the MAP-MDS estimator[37] by showing the shortest path algorithm faithfully returns the Euclidean distance between unconnected nodes when the point set is regular. In parallel, [25] obtained almost the same result, and they sent the surrogate produced by MAP-MDS into the OptSpace[33] for refinement. Numerically, it was demonstrated that the phase transition of OptSpace matches that of the theoretical bound on MAP-MDS. [16, Thm. 1.2] provided bounds on similar refinement step, yet they treat the EDM-dependent samples as pure additive noise and only address the Bernoulli erasures.
@@
The original motivation for raising Q1 comes from Tasissa and Lai[20], who showed, for the first time that a natural parallelization of the Nuclear Norm Minimization (NNM)[38][39], called trace minimization (the trace regularized version of the Semidefinite Relaxation (SDR) of (5)), can solve EDMC given $\mathcal{O}(n\nu r\log^{2}n)$ uniform distance samples taking at random, where $\nu$ is certain coherence parameter defined w.r.t. the EDM basis $\boldsymbol{\omega}_{\boldsymbol{\alpha}}$ [^2] The form of $\boldsymbol{\omega}_{\boldsymbol{\alpha}}$ will be introduced shortly in Section II. We note that $\nu=\mathcal{O}(r\mu)$, where $\mu$ is the standard coherence parameter defined w.r.t. $\mathbf{e}_{i}\mathbf{e}_{j}^{T}$, c.f., [38, Def. 1.2].. They also demonstrated the generalizability of this specific modification to the Golfing Scheme[40] by analyzing matrix completion problem under arbitrary set of non-orthonormal and non-sub-Gaussian basis, provided the extremal eigenvalues of the basis correlation matrix are well understand[41]. Subsequent work[42] characterized the eigen-spectrum of the correlation matrix $\mathbf{H}_{\boldsymbol{\omega}}$ of $\boldsymbol{\omega}_{\boldsymbol{\alpha}}$ basis, and revealed that $\mathbf{H}_{\boldsymbol{\omega}}$ has only three types of distinct eigenvalues: $2n,\,n$, and $2$. Such near-singular spectral structure does not significantly hinder the convergence of the Golfing Scheme, as the original inexact dual certificate requires precision on $\mathcal{O}(1/n)$ for the residual term $\|\mathcal{P}_{\mathbb{T}}\mathbf{Y}-\mathrm{sgn}(\mathbf{X}^{\star})\|$[32, Prop. 2], while the modified one for EDMC appears like $\mathcal{O}(\frac{\lambda_{\min}(\mathbf{H}_{\boldsymbol{\omega}})}{n\lambda_{\max}(\mathbf{H}_{\boldsymbol{\omega}})})=\mathcal{O}(1/n^{2})$[41, Thm. 2]. Since the Golfing Scheme converges exponentially fast, the extra iteration cost for improving $1/n$ to $1/n^{2}$ is negligible. Yet this makes the design and analysis of non-convex methods considerably more obscure, which, in its most straightforward form, aims to solve a fourth-order polynomial as follows
@@
$[\mathcal{P}_{\Omega}(\mathbf{D})]_{ij}$ =$=\begin{cases}\mathbf{D}_{ij},\,\,&\text{if}\,\,(i,j)\in\Omega\\ 0,\,\,&\text{else}\end{cases}.$
where $\Omega$ is the sample set, $\boldsymbol{\alpha}=(i,j)$ is a tuple and the distance between $(i,j)$ points, $(i,j)\in\Omega$, is known. $\mathbf{D}^{\star}_{\boldsymbol{\alpha}}=\langle\mathbf{D}^{\star},\mathbf{e}_{i}\mathbf{e}_{j}^{T}\rangle=\mathbf{D}^{\star}_{ij}$ denote the sampled ground truth EDM, where $\mathbf{e}_{i}$ stands for the $i$-th canonical Euclidean basis in $\mathbb{R}^{n}$. $\mathbf{P}$ stands for the point set to be recovered. $\mathcal{P}_{\Omega}g(\cdot)$ can be viewed as the matrix wrap up of the expression $\sum_{\boldsymbol{\alpha}\in\Omega}\langle\cdot,\boldsymbol{\omega}_{\boldsymbol{\alpha}}\rangle\mathbf{e}_{i}\mathbf{e}_{j}^{T}$ for the moment, and the precise form of $g$ will be shortly defined below. Eq. (1) bears great resemblance to the Wirtinger Flow[43] used in phase retrieval, Hankel lift for spectral compressive sensing[44][45], $l_{2,\infty}$ regularized free non-convex matrix completion[46][47], and so on [^3] Historically speaking, (1) is called the s-stress function[48], interested reader is referred to our previous work for a short survey[22].. However, the large condition number of $\mathbf{H}_{\boldsymbol{\omega}}$ prevents us from pursuing theoretical contraction speed when optimizing (1), as discussed in [22, Sec. III-C][23, Sec. 6], even though (1) does produce an attractive basin given enough Bernoulli samples, the basin is too small when measured by $l_{2,\infty}$ norm.
@@
Another interesting open question concerning (1) is the analytical justification for the observed absence of its spurious local minima, as summarized and empirically reported in[51, Ch. 3.5]. This phenomenon is particularly striking: drawing from the global landscape analysis of Zhu et al.[52, Thm. 1], one might expect that ill-conditioning of the Hessian associated with the un-factorized cost function would preclude the existence of a benign optimization landscape. Moreover, due to the lost of a smooth generalization of the random graph lemma[33, Lemma 7.1] to the EDMC settings as mentioned above, Ge et al.’s framework[53][54], i.e., by introducing penalty term on the $l_{2,\infty}$ norm of $\mathbf{P}$ into (1) to amend the curvature of landscape, cannot be immediately adopted. As we will elaborate further, it also remains an open question whether the standard incoherence condition adequately captures the class of EDMs that are efficiently completable in polynomial time [^4] Please see [21, Thm. 4] for more discussion..
@@
where $\mathcal{S}^{n,r}_{+}$ denotes the Positive Semidefinite (PSD) fix-rank embedding matrix manifold[55]. The major drawback is that one cannot compute $\nabla F(\mathbf{G})$ but only a part of it, which they called pseudo gradient ^$\hat{\nabla}F(\mathbf{G}):=\frac{2}{p}\mathcal{R}_{\Omega}(\mathbf{G}-\mathbf{G}^{\star})$. Certain trade-off arises here between maintaining self-adjointness and ensuring algorithmic explainability: since $\mathbf{H}_{\boldsymbol{\omega}}^{-1}$ (the inverse of $\mathbf{H}_{\boldsymbol{\omega}}$) is not diagonal, in general we would like to evade from constructing $\mathcal{R}_{\Omega}^{*}\mathcal{R}_{\Omega}$ [^5] Finding uniform bounds on $\mathcal{R}_{\Omega}^{*}\mathcal{R}_{\Omega}$, e.g., analogy of Lemma C.3, C.5 is challenging. While treating it as a sub-Gaussian quadratic form and applying the “suprema of chaos” inequality[56] as in [57][58] might yield a similar tangent space eigenvalue lower bound to[20, Lemma 10], the upper bound for replacing Lemma C.5 remains too loose., while using $\mathcal{R}_{\Omega}$ alone compromises the least-squares structure. Things tend to be more subtle when one inquires about numerical performance. This Riemannian pseudo gradient descent does not exhibit strong convergence in the sample-limited regime when compared to (1), despite its theoretical error contraction rate resembling that of the classic Riemannian Gradient Descent for Low-rank Matrix Completion (LRMC)[59].
I-B Major Contributions
By introducing Burer-Monteiro factorization into (3), we have

### v2

@@ Abstract
This paper proposes and analyzes a gradient-type algorithm based on Burer-Monteiro factorization, called the Asymmetric Projected Gradient Descent (APGD), for reconstructing the point set configuration from partial Euclidean distance measurements, known as the Euclidean Distance Matrix Completion (EDMC) problem. By paralleling the incoherence matrix completion framework, we show for the first time that global convergence guarantee with exact recovery of this routine can be established given $\mathcal{O}(\mu^{2}r^{3}\kappa^{2}n\log n)$ Bernoulli random observations without any sample splitting. Unlike leveraging the tangent space Restricted Isometry Property (RIP) and local curvature of the low-rank embedding manifold in some very recent works, our proof provides extra upper bounds that act as analogies of the random graph lemma under EDMC setting. The APGD works surprisingly well and numerical experiments demonstrate exact linear convergence behavior in rich-sample regions yet deteriorates rapidly when compared with the performance obtained by optimizing the s-stress function, i.e., the standard but unexplained non-convex approach for EDMC, if the sample size is limited. While virtually matching our theoretical prediction, this unusual phenomenon might indicate that: (i) the power of implicit regularization is weakened when specified in the APGD case; (ii) the stabilization of such new gradient direction requires substantially more samples than the information-theoretic limit would suggest.
@@ I Introduction
Given $n$ points $\{\mathbf{p}_{i}^{\star}\}_{i=1}^{n}$ embedded in $\mathbb{R}^{r}$, $r=2,3$, calculating their inter-point distances is a trivial task. Forwarding the perfect, complete n(n1)/2n(n-1)/2 distances back to the point set configuration is, meanwhile well-known as the classic Multi-dimensional Scaling (cMDS)[2, Sec. 3.1.1][3][4, Thm. 2]. However, significant effort has been devoted to handle the case when part of the distance measurements are missing, known as the Euclidean Distance Matrix Completion (EDMC) problem[5][4]. This is not only due to its ubiquitous presence in, e.g., computational chemistry[6][7][8][9], operating wireless sensor networks[10][11][12] and articulated robots[5, Sec. 4.3.2][13], channel estimation[14], indoor SLAM[15], ultrasound calibration[16], but also the fact that this inverse problem is NP-hard to solve in general[17, Sec. 3.4]. To circumvent the NP-hardness, one may only consider a selected region of the ground truth (generic[18][19], incoherence[20][21]), enforcing certain connectivity of the underlying sample graph (Erdös-Rényi[22][23], random geometric graph[12][24][25][26]), and posing constrains on certain geometry of the point set ($r$-unique localizable[27], trilateration[28], universal rigidity[29]). Therefore, performance guarantee results in this field can be roughly divided into two classes, i.e., combinatorial[27][29][30] and probabilistic. While the latter can be further categorized into the traditional yet effective view from random geometry[12][24][25][26], and the recent revisit of EDMC via incoherent matrix completion lens[20][23][31][22]. Method proposed in this manuscript belongs to the last class. In short, we analyze an algorithm analogous to the Iterative Fast Hard Thresholding for EDMC (IFHT-EDMC) that appeared in a very recent work [23], under standard incoherence assumption[32] and Bernoulli sampling scheme, from the quotient geometry perspective.
I-A Preliminaries and Motivations
A Euclidean Distance Matrix (EDM), $\mathbf{D}^{\star}\in\mathbb{R}^{n\times n}$, is referred to be a symmetric hollow diagonal matrix whose (i,j)(i,j) [^1] We will frequently use boldface Greek letter $\boldsymbol{\alpha}$ or $\boldsymbol{\beta}$ to denote an index tuple $(i,j)\in\mathbb{I}$, as inherited from[20]. entity denotes the squared distance between $(i,j)$ points
$\mathbf{D}^{\star}_{ij}:=\|\mathbf{p}_{i}^{\star}-\mathbf{p}_{j}^{\star}\|_{2}^{2}=\langle\mathbf{P}^{\star}\mathbf{P}^{\star T},\boldsymbol{\omega}_{\boldsymbol{\alpha}}\rangle=\langle\mathbf{G}^{\star},\boldsymbol{\omega}_{\boldsymbol{\alpha}}\rangle,$
where $\langle\mathbf{A},\mathbf{B}\rangle:=\mathrm{tr}(\mathbf{A}^{T}\mathbf{B})$ is the Frobenius inner product. $\mathbf{P}^{\star}:=[\mathbf{p}_{1}^{\star},...,\mathbf{p}_{n}^{\star}]^{T}\in\mathbb{R}^{n\times r}$ and $\mathbf{G}^{\star}:=\mathbf{P}^{\star}\mathbf{P}^{\star T}$ are the so-called point set configuration and Gram matrix [^2] The point set configuration is assumed to have full column rank $r$., respectively. Let $\mathbf{e}_{i}$ denote the $i$-th canonical Euclidean basis in $\mathbb{R}^{n}$, then $\boldsymbol{\omega}_{\boldsymbol{\alpha}}:=(\mathbf{e}_{i}-\mathbf{e}_{j})(\mathbf{e}_{i}-\mathbf{e}_{j})^{T}$ stands for the (primal) EDM basis. The EDMC problem considered here aims at recovering an EDM from its partial and element-wise measurements, which is usually referred to be the following non-convex QCQP (Quadratically Constrained Quadratic Programming)
$\mathrm{find}$ $\mathbf{G}$ (1)
$\mathrm{s.t.}$ $\langle\mathbf{G},\boldsymbol{\omega}_{\boldsymbol{\alpha}}\rangle=\mathbf{D}^{\star}_{\boldsymbol{\alpha}},\,\boldsymbol{\alpha}=(i,j)\in\Omega,$
$\mathbf{G1}=\mathbf{0},\,\mathbf{G}\succcurlyeq\mathbf{0},\,\mathrm{rank}(\mathbf{G})=r,$
where $\mathbf{D}^{\star}_{\boldsymbol{\alpha}}:=\mathbf{D}^{\star}_{ij}$. Superscript $(\cdot)^{\star}$ is used to emphasize the fixed, ground truth that independent with the sample set $\Omega\subset\mathbb{I}:=\{(i,j):1\leq i<j\leq n\}$ [^3] Since the EDM is a symmetric and hollow matrix, we can omit its diagonal and only sample the upper triangle sub-matrix, which is referred to as the symmetric sampling scheme later., and the distances between $(i,j)$ points, $(i,j)\in\Omega$, are known. Since rigid motions in $\mathbb{R}^{r}$ preserve the relative distance between points, the self-centered constraint $\mathbf{G1}=\mathbf{0}$, is usually adopted to remove the translation ambiguity[4, Eq. (29)], where $\mathbf{1}$, $\mathbf{0}$ are the vectors of all ones, all zeros, respectively.
If the underlying ground truth framework (point set configuration together with the sample graph $\Omega$) is universally rigid, then the rank constraint can be removed at no cost, resulting in streamlined processing on (1) with modern convex solvers[29]. However, two unresolved issues remain and can not be compromised[22][23][31][33]: (i) non-degenerate trilateration graphs[28] are known to be universally rigid[29, Coro. 1], yet this relationship is nowhere inclusive. Moreover, to the best of our knowledge, there are no effective algorithms to test the universal rigidity except solving the Biswas-Ye Semidefinite Relaxation (SDR)[27]; (ii) Semidefinite Programming (SDP) based approaches convexify the matrix dimension to $n\times n$, causing extra computational burden when $n\gg r$. Skeptical readers may argue that (1) can be tackled by convex-lifting (over-parameterization) to meet the SDR rank lower bound[34], combined with sophisticated manifold optimization strategies, e.g.,[35, Sec. 4.4]. Yet the success of such routine does not explain practical observations: (i) usually small over-parameterization suffices for non-convex algorithms to evade saddle points[22, Sec. III-D] and local minima[33, Sec. 7]; (ii) the recommended numerical method to obtain a successful recovery remains low-rank inductive regularization[9, Sec. 3.1][20, Sec. IV]. Therefore, we are interested in answering the following basic question [^4] Please notice that low-rankness alone can not fully characterize an EDM[4], and direct adaptation of Burer-Monteiro factorization or nuclear norm minimization to an EDM without leveraging rank-one quadratic sensing structure w.r.t. $\boldsymbol{\omega}_{\boldsymbol{\alpha}}$ basis in (1) leads to poor performance[22, Sec. I-B].
@@ I-A Preliminaries and Motivations
Q1: Does the theoretical success of the classical two-step non-convex factorization type incoherent Low-rank Matrix Completion (LRMC) algorithms [36][37][38][39] generalize to the EDMC setting?
As depicted in our previous work[22], a twofold dilemma arises when addressing Q1. First, practical sampling schemes in real-world applications of the EDMC problem often deviate from the ideal Bernoulli model. Second, the correlation matrix of the $\boldsymbol{\omega}_{\boldsymbol{\alpha}}$ basis (denoted as $\mathbf{H}_{\boldsymbol{\omega}}$, c.f. (2)) is non-diagonal, causing simple re-weighting scheme as in Hankel Matrix Completion (HMC)[40][41] no longer valid. This work addresses the second challenge by providing the first non-convex routine to solve EDMC with global convergence guarantees, without requiring sample splitting, and our approach leverages different proof techniques compared to[22][23][31][42]. We now discuss these and highlight the difference in greater detail.
I-B Related Work
@@
While the Bernoulli sample model used in standard LRMC tasks does not have strong real-world application scenarios when specified to EDMC, we are not aware of any work that analyzed the non-convex gradient refinement stage under a more practical sample model that depends on the magnitudes of EDM entities, i.e., the so-called unit ball rule, under the random geometry setup [^5] Transition bound obtained from random geometry graph is often sharp up to constant, yet it also requires ideal distribution assumptions on the point set.. Javanmard and Montanari developed the stability notion of a rigid graph[12, Lemma 4.1, 4.2] and use that to show the perturbation on convex solution can be bounded. [26][24] analyzed the MAP-MDS estimator[43] by showing the shortest path algorithm faithfully returns the Euclidean distance between unconnected nodes when the point set is regular. In parallel, [25] obtained almost the same result, and they sent the surrogate produced by MAP-MDS into the OptSpace[36] for refinement. Numerically, it was demonstrated that the phase transition of OptSpace matches that of the theoretical bound on MAP-MDS. [16, Thm. 1.2] provided bounds on similar refinement step, yet they treat the EDM-dependent samples as pure additive noise and only address the Bernoulli erasures.
@@
The original motivation for raising Q1 comes from Tasissa and Lai[20], who showed, for the first time that a natural parallelization of the Nuclear Norm Minimization (NNM)[44][45], called trace minimization (the trace regularized version of the SDR of (1)), can solve EDMC given $\mathcal{O}(n\nu r\log^{2}n)$ uniform distance samples taking at random, where $\nu$ is certain coherence parameter defined w.r.t. the EDM basis $\boldsymbol{\omega}_{\boldsymbol{\alpha}}$ [^6] In general $\nu=\mathcal{O}(r\mu)$, where $\mu$ is the standard coherence parameter defined w.r.t. $\mathbf{e}_{i}\mathbf{e}_{j}^{T}$, c.f., [44, Def. 1.2].. They also demonstrated the generalizability of this specific modification to the Golfing Scheme[46] by analyzing matrix completion problem under an arbitrary set of non-orthonormal and non-sub-Gaussian basis, provided the extremal eigenvalues of the basis correlation matrix are well understood[47]. Subsequent work [48] characterized the eigen-spectrum of the EDM basis correlation matrix $\mathbf{H}_{\boldsymbol{\omega}}$, which is defined as
$\mathbf{H}_{\boldsymbol{\omega}}:=\mathbf{WW}^{T}\in\mathbb{R}^{L\times L},\,L:=n(n-1)/2=|\mathbb{I}|,$ (2)
$\mathbf{W}:=[\mathrm{vec}(\boldsymbol{\omega}_{(1,2)}),...,\mathrm{vec}(\boldsymbol{\omega}_{(n-1,n)})]^{T}\in\mathbb{R}^{L\times n^{2}},$
where $\mathrm{vec}(\cdot)$ stands for matrix vectorization. [48, Coro. 2.2] revealed that $\mathbf{H}_{\boldsymbol{\omega}}$ has only three types of distinct eigenvalues: 2n,n2n,\,n, and $2$. Such near-singular spectral structure does not significantly hinder the convergence of the Golfing Scheme, as the original inexact dual certificate requires precision on $\mathcal{O}(1/n)$ for the residual term[32, Prop. 2-(2a)], while the modified one for EDMC appears like $\mathcal{O}(\frac{\lambda_{\min}(\mathbf{H}_{\boldsymbol{\omega}})}{n\lambda_{\max}(\mathbf{H}_{\boldsymbol{\omega}})})=\mathcal{O}(1/n^{2})$[47, Thm. 2]. Since the Golfing Scheme converges exponentially fast, the extra iteration cost for improving $1/n$ to $1/n^{2}$ is negligible. Yet this makes the design and analysis of non-convex methods considerably more obscure, which, in its most straightforward form, aims to solve the s-stress cost function defined below
@@
$[\mathcal{P}_{\Omega}(\mathbf{D})]_{ij}$ $=\begin{cases}\mathbf{D}_{ij},\,\,&\text{if}\,\,(i,j)\in\Omega\\ 0,\,\,&\text{else}\end{cases}.$
Where standard notation $\mathcal{P}_{\Omega}$ is adopted to define the projection onto sampled entities, and linear operator $g:\mathcal{S}(n)\to\mathcal{S}(n)$ maps Gram matrix $\mathbf{G}=\mathbf{PP}^{T}$ to the EDM generated by $\mathbf{P}$. Here, $\mathcal{S}(n)$ denotes the set of symmetric matrix in $\mathbb{R}^{n\times n}$, see Section II for formal definitions of $g$. Eq. (3) bears great resemblance to the Wirtinger Flow[49] used in phase retrieval, Hankel lift for spectral compressive sensing[41][50], $l_{2,\infty}$ regularized free non-convex matrix completion[51][52], and so on. However, the large condition number of $\mathbf{H}_{\boldsymbol{\omega}}$ prevents one from pursuing theoretical contraction speed when optimizing (3), as discussed in [22, Sec. III-C][23, Sec. 6], even though (3) does produce an attractive basin given enough Bernoulli samples, the basin is too small when measured by $l_{2,\infty}$ norm.
@@
TABLE I: Comparison with recent works based on dual basis expansion
Algorithm Type Incoherence Type Claimed Sample Complexity Global Recovery Sample Splitting Riemannian Geometry
Trace Minimization[20] Strong $\mathcal{O}(nr\log^{2}n)$[20, Thm. 5] N.A. N.A.
Iterative Fast Hard Thresholding[23] Strong $\mathcal{O}(nr^{2}\log^{2}n)$[23, Thm. 5.9] Embedding
Iterative Re-weighting Least Square[31] Strong $\mathcal{O}(nr\log n)$[31, Thm. 4.3] $\times$ $\times$ N.A.
Projected Gradient Descent (This work) Standard $\mathcal{O}(nr^{3}\log n)$ $\times$ Quotient
Another long-standing yet recently advanced open question concerning (3) is the analytical justification for the observed absence of its spurious local minima, as summarized and empirically reported in[55, Ch. 3.5]. When the observation is complete and noiseless, this conjecture was recently disproven by Criscitiello et al.[33, Sec. 6]. They construct explicit spurious configurations and provide sufficient condition for s-stress to admit strict second-order local minima. An interesting yet subtle connection between their work and ours lies in[33, Thm. 5.1]. Loosely speaking, they show that if
$(k+2)\sigma_{r}^{\star}>4n\max_{i\in[n]}\|\mathbf{e}_{i}^{T}\mathbf{P}^{\star}\|_{2}^{2},\,\Omega=\mathbb{I},$ (4)
holds, then any over-parameterized second-order critical $\mathbf{P}\in\mathbb{R}^{n\times k}$, with $k>r$, is a global optimum of (3), where $\sigma_{r}^{\star}$ is the smallest singular value of $\mathbf{G}^{\star}$. Though sharing remarkable resemblance with the incoherence assumption[44, Def. 1.2], it is straightforward to check (4) can only hold in the regime $k>r$. And as we will elaborate further, it also remains an open question whether the standard incoherence condition adequately captures the class of EDMs that are efficiently completable in polynomial time [^7] Please see [21, Thm. 4] for more discussion..
@@
$\nabla F(\mathbf{G})=\frac{1}{p}\mathcal{R}_{\Omega}(\mathbf{G}-\mathbf{G}^{\star})+\frac{1}{p}\mathcal{R}_{\Omega}^{*}(\mathbf{G}-\mathbf{G}^{\star}),$
where $\mathcal{S}^{n,r}_{+}$ denotes the Positive Semidefinite (PSD) rank-$r$ embedding matrix manifold[56]. The major drawback is that one cannot compute $\nabla F(\mathbf{G})$ directly, but only its first term, which they called pseudo gradient (after proper rescaling) $\hat{\nabla}F(\mathbf{G}):=\frac{2}{p}\mathcal{R}_{\Omega}(\mathbf{G}-\mathbf{G}^{\star})$, since the second term requires precise knowledge of $\mathbf{G}^{\star}$ but not just sampled distances. Certain trade-off arises here between maintaining self-adjointness and ensuring algorithmic explainability: since $\mathbf{H}_{\boldsymbol{\omega}}^{-1}$ (the inverse of $\mathbf{H}_{\boldsymbol{\omega}}$) is not diagonal, in general we would like to evade from constructing $\mathcal{R}_{\Omega}^{*}\mathcal{R}_{\Omega}$ [^8] Finding uniform bounds on $\mathcal{R}_{\Omega}^{*}\mathcal{R}_{\Omega}$, e.g., analogy of Lemma C.3, C.5 is challenging, we will come back to this with greater detail in Section II-C., while using $\mathcal{R}_{\Omega}$ alone compromises the least-squares structure. The situation becomes more subtle when one inquires about numerical performance. This Riemannian pseudo gradient descent does not exhibit strong convergence in the sample-limited regime when compared to (3), despite its theoretical error contraction rate resembling that of the classic Riemannian gradient descent for LRMC[57].
I-C Major Contributions and Organization
By introducing Burer-Monteiro factorization into (6), and recalculating the pseudo gradient, we have